# DBellQuant: Breaking the Bell with Double-Bell Transformation for LLMs Post Training Binarization

### v1

@@ A.2 Broader Impacts
A.4 Analysis of the reasons doubel-bell distribution more suitable for binarization compared to a single-bell distribution.
@@
A.4.1 Scaling Up Small Values Reduces Quantization Error
Consider the scenario where the input vector is $X=[2,2,2,\dots,2]_{n}$. Assume the weights in a single channel, $W$, are given by $[\alpha_{1},\alpha_{2},...,\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $[\alpha_{i}\in U_{1}]$ is the set of values with very large absolute magnitudes, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $[\beta_{i}\in U_{2}]$ are values with magnitudes close to zero, satisfying $\sum_{i=1}^{n-k}\beta_{i}=0$. $W=U_{1}\cup U_{2}$. This structure is common in practice, as weight distributions in neural networks often exhibit a few dominant values and many small ones. As we observe, the distribution of weights along the channel dimension is mostly not symmetric around zero. Instead, it tends to be biased, with the majority leaning either towards positive or negative values. Consequently, the extreme values are predominantly either entirely positive or entirely negative, so we assume $\alpha_{i}>0$.
@@ A.4.1 Scaling Up Small Values Reduces Quantization Error
Given that $[\beta_{1},\beta_{2},\dots,\beta_{n-1}]$ are all very small in magnitude compared to $\frac{\alpha}{n}$, we can simplify the above as:
$\begin{split}AbsMean&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-\beta_{1})+(\frac{A}{n}-\beta_{2})+\dots+(\frac{A}{n}-\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{2kA}{n})}{n}\end{split}$ (10)
@@
$-AbsMean+M=-\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (12)
@@
Let us now analyze the quantization error. The original output is $2\alpha$. For convenience, let $N=(n-k)(-AbsMean+M)$. The quantized output is a sum of the dequantized values for all weights.
@@
A.4.2 Scaling Down Large Values Reduces Quantization Error
Now, let us consider the scenario where we scale down the large-magnitude weight. Let the input $X=[1,1,1,\dots,1]_{n}$, and the weights $W=[m\alpha_{1},m\alpha_{2},...,m\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ are large values, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\alpha_{i}>0$, $[\beta_{1},\dots,\beta_{n-k}]$ are small, and $\sum_{i=1}^{n-k}\beta_{i}=0$.
@@ A.4.2 Scaling Down Large Values Reduces Quantization Error
$-AbsMean+M=-\frac{2m(A-\frac{A}{n})}{n}+\frac{mA}{n}=\frac{-mA+\frac{2mkA}{n}}{n}$ (26)
@@
$-AbsMean_{new}+M_{new}=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{-A+\frac{2kA}{n}}{n}$ (32)
@@
Because $A>0$, for $n>100$ ,$\alpha>0$ and $k\ll n$ it holds that:
@@
A.5 Perplexity on C4 dataset
[…]
Full Precision 16 16.07 14.34 12.71 7.34 07.26 06.73 5.71
RTN 16 $9$.56 23492.89 9617.07 194607.78 115058.76 46250.21 314504.09
GPTQ 16 6364.65 6703.36 5576.82 186229.5 67954.04 19303.51 13036.32
PB-LLM 16 168.12 222.15 104.78 76.63 80.69 184.67 NAN
BiLLM 16 64.14 44.77 42.13 46.96 39.38 25.87 17.30
[…]
BiLLM 8 74.56 61.99 40.91 47.13 40.91 21.45 17.72
[…]
BiLLM 6 7348 13445.21 63.41 61.65 63.41 37.66 19.43
[…]
A.6 Visualization of distribution of activation before and after DBellQuant

### v2

@@ A.18 Algorithm
A.20 Analysis of the reasons double-bell distribution more suitable for binarization compared to a single-bell distribution.
@@
A.20.1 Scaling Up Small Values Reduces Quantization Error
Consider the scenario where the input vector is $X=[2,2,2,\dots,2]_{n}$. Assume the weights in a single channel, $W$, are given by $\alpha_{1},\alpha_{2},...,\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{i}\in U_{1}]$ is the set of values with very large absolute magnitudes, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\beta_{i}\in U_{2}]$ are values with magnitudes close to zero, satisfying $\sum_{i=1}^{n-k}\beta_{i}=0$. $W=U_{1}\cup U_{2}$. This structure is common in practice, as weight distributions in neural networks often exhibit a few dominant values and many small ones. As we observe, the distribution of weights along the channel dimension is mostly not symmetric around zero. Instead, it tends to be biased, with the majority leaning either towards positive or negative values. Consequently, the extreme values are predominantly either entirely positive or entirely negative, so we assume $\alpha_{i}>0$.
@@ A.20.1 Scaling Up Small Values Reduces Quantization Error
Given that $\beta_{1},\beta_{2},\dots,\beta_{n-1}]$ are all very small in magnitude compared to $\frac{\alpha}{n}$, we can simplify the above as:
$\begin{split}AbsMean&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-\beta_{1})+(\frac{A}{n}-\beta_{2})+\dots+(\frac{A}{n}-\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{kA}{n})}{n}\end{split}$ (10)
@@
$AbsMean+M=-\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (12)
@@
Let us now analyze the quantization error. The original output is $2\alpha$. For convenience, let N=(nk)(AbsMean+M)N=(n-k)(-AbsMean+M). The quantized output is a sum of the dequantized values for all weights.
@@
A.20.2 Scaling Down Large Values Reduces Quantization Error
Now, let us consider the scenario where we scale down the large-magnitude weight. Let the input $X=[1,1,1,\dots,1]_{n}$, and the weights $W=[m\alpha_{1},m\alpha_{2},...,m\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ are large values, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\alpha_{i}>0$, $\beta_{1},\dots,\beta_{n-k}]$ are small, and $\sum_{i=1}^{n-k}\beta_{i}=0$.
@@ A.20.2 Scaling Down Large Values Reduces Quantization Error
$AbsMean+M=-\frac{2m(A-\frac{A}{n})}{n}+\frac{mA}{n}=\frac{-mA+\frac{2mkA}{n}}{n}$ (26)
@@
$AbsMean_{new}+M_{new}=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{-A+\frac{2kA}{n}}{n}$ (32)
@@
Because $A>0$, for $n>100$ $\alpha>0$ and $k\ll n$ it holds that: